# Look Around and Find Out: OOD Detection with Relative Angles

### v1

@@ Abstract
Deep learning systems deployed in real-world applications often encounter data that is different from their in-distribution (ID). A reliable system should ideally abstain from making decisions in this out-of-distribution (OOD) setting. Existing state-of-the-art methods primarily focus on feature distances, such as k-th nearest neighbors and distances to decision boundaries, either overlooking or ineffectively using in-distribution statistics. In this work, we propose a novel angle-based metric for OOD detection that is computed relative to the in-distribution structure. We demonstrate that the angles between feature representations and decision boundaries, viewed from the mean of in-distribution features, serve as an effective discriminative factor between ID and OOD data. Our method achieves state-of-the-art performance on CIFAR-10 and ImageNet benchmarks, reducing FPR95 by $0.88\%$ and $7.74\%$ respectively. Our score function is compatible with existing feature space regularization techniques, enhancing performance. Additionally, its scale-invariance property enables creating an ensemble of models for OOD detection via simple score summation. Code available at https://github.com/berkerdemirel/LAFO-Look-Around-and-Find-Out
@@ 1 Introduction
A trustworthy deep learning system should not only produce accurate predictions, but also recognize when it is processing an unknown sample. The ability to identify when a sample deviates from the expected distribution, and potentially rejecting it, plays a crucial role especially in safety-critical applications, such as medical diagnosis  [11], driverless cars  [3] and surveillance systems  [7]. The out-of-distribution (OOD) detection problem addresses the challenge of distinguishing between in-distribution (ID) and OOD data – essentially, drawing a line between what the system knows and what it does not.
@@
Various approaches have been proposed for OOD detection, mainly falling into two categories: (i) methods that suggest model regularization during training  [22, 16, 28], and (ii) post-hoc methods, which leverage a pre-trained model to determine if a sample is OOD by designing appropriate score functions  [35, 15, 39]. Post-hoc methods are more advantageous for their efficiency and flexibility, as they can be applied to arbitrary pre-trained models without retraining. These approaches are often categorized based on the domain of their score functions, i.e., at which representational abstraction level they assess if a sample is OOD or not. Earlier techniques focus on measuring the model confidence in the logits space  [15, 26], while the recent works employ distance-based scores  [39, 36] defined in the model feature space. While logit-based methods suffer from the overconfident predictions of neural networks  [29, 21, 13], the recent success of distance-based techniques highlights that the relationships in the latent space can provide a richer analysis.
@@
A natural approach to feature representations is by checking their proximity to the decision boundaries  [25]. Conceptually, this can be related to identifying hard-to-learn examples in data-efficient learning  [18, 4]. OOD samples can be viewed as hard-to-learn since they do not share the same label distribution as ID data. The success of this approach has been directly showed in fDBD score from  [25]. However, our derivations revealed that the regularization term they use to incorporate ID statistics introduces an additional term that does not correlate with ID/OOD separation, ultimately hindering their performance.
@@
In this work, we present Look Around and Find Out (LAFO), a novel approach that exploits the relationship between feature representations and classifier decision boundaries, in the context of the mean statistics of ID features. Unlike the earlier techniques, LAFO introduces a new angle-based measure that calculates the angles between the feature representations and their projection onto the decision boundaries, relative to the the mean statistics of ID features. Changing reference frame to the mean of ID features adds another layer of discriminatory information to the score, as it naturally incorporates the ID statistics to the distance notion, exploiting the disparity between ID and OOD statistics. Moreover, the scale-invariant nature of angle-based representations, as similarly observed in [32], allows us to aggregate the confidence scores from multiple pre-trained models simply by summing their LAFO scores. This enables to have a score that can be single model based or extended to ensemble of models. In summary, our key contributions include:
@@
  • We present a novel technique for OOD detection, which computes the angles between the feature representation and its projection to the decision boundaries, relative to the mean of ID-features
  • Our score is model agnostic, hyperparameter-free and efficient, scaling linearly with the number of ID-classes. Therefore, it can flexibly be combined with various architectures without the need of additional tuning.
  • We demonstrate the state-of-the-art performance of LAFO on widely used CIFAR-10 and ImageNet OOD benchmarks. Specifically, LAFO achieves a $7.74\%$ reduction in FPR95 score compared to the best previous distance-based method on a large-scale ImageNet OOD benchmark.
@@
  • The scale-invariant property of LAFO allows for straightforward aggregation of confidence scores from multiple pre-trained models, improving ensemble performance. Our experiments show that the ensemble with LAFO reduces the FPR95 by $2.51\%$ on the ImageNet OOD benchmark compared to the best single model performance.
@@ 2 Related Work
Early methods addressing the OOD detection problem  [2, 16, 31, 50] utilize additional datasets to represent out-of-distribution data, training models with both positive and negative samples. This approach assumes a specific nature of OOD data, potentially limiting its effectiveness when encountering OOD samples that deviate from this assumption during inference.  [27] designed a network architecture to measure distributional uncertainty using Dirichlet Prior Networks. In  [12]’s work, they provide another architecture with an additional reject option to abstain from answering. Their selection model incorporates a hyperparameter, the coverage rate for ID, to control the percentage of ID samples classified.  [22, 30, 10], focused on synthesizing outliers rather than relying on auxilary datasets to improve the generalizability of the detection method. On the other hand,  [28, 41, 44] argued that overconfident predictions of the networks on OOD data are the problem to be mitigated. For example,  [41] puts an additional gradient penalty to limit the confidence of the network. Whereas,  [44] tackled the same problem by enforcing a constant logit vector norm during training. Although it is natural to impose structures during training for better separability of ID and OOD, these methods face the trade-off between OOD separability and model performance. Moreover, such approaches lack the flexibility of post-hoc score functions, as they necessitate model retraining—a process that can be both time-consuming and computationally expensive.
Score Functions
@@
Recently, developing score functions for pretrained models on ID data has gained attention due to its ease of implementation and flexibility. These methods typically either couple feature representations with distance metrics, or measure a model’s confidence using its logits. Beyond canonical works such as Maximum Softmax Probability  [15], ODIN score  [24], Energy score  [26], Mahalanobis score  [23], Virtual Logit Matching  [43], we observed many advancements in post-hoc score design. For example, the activation shaping algorithms such as ASH  [8], Scale  [47], and ReAct  [38], apply activation truncations to feature representations, reducing model’s confidence for OOD data. These methods can be used in conjunction with LAFO improving the performance. Following a similar intuition, DICE  [37] applies weight sparsification to limit the model confidence by using the more salient weights for output. GradOrth  [1] offers a gradient-based perspective, projecting representations to a lower dimensional subspace based on gradient norms. Recent distance-based methods KNN  [39] and FDBD  [25] successfully utilized the feature representations from networks trained with supervised contrastive loss. KNN assigns a score to a sample based on the kth nearest neighbor in ID training data. On the other hand, FDBD assigns a score to a sample based on its estimate of the distance between the feature representation and the decision boundaries.
@@
Our work falls into the score function category, serving as a plug-in for any pre-trained model on ID data. LAFO combines feature space and logit space methods by utilizing the relative angle between the feature representation and its projections to the decision boundaries. Among the existing works, the closest approach to our method is  [25], which uses a lower bound estimate to the decision boundaries. However, the regularization term they introduced to equalize deviations from the in-distribution mean inadvertently includes a term in their equation that is uncorrelated with being out- or in-distribution and can change spuriously, impeding performance. In contrast, we provide a simple, hyperparameter-free score function that effectively incorporates in-distribution context and maintains scale invariance, all without extra regularization terms.
@@
Figure 1: Geometric visualization of LAFO for in-distribution (left) and out-of-distribution (right) cases. LAFO focuses on the angular distance between the feature representation and the decision boundary, from the perspective of the in-distribution mean. The angle $\theta$ serves as the distinguishing factor between ID and OOD samples, with $\theta_{\text{ID}}>\theta_{\text{OOD}}$.
@@ 3 Method
3.1 Problem Setting
We consider a supervised classification setting with input space $\mathcal{X}$ and label space $\mathcal{Y}$, following the literature  [51]. Given a model $f:\mathcal{X}\rightarrow\mathbb{R}^{|\mathcal{Y}|}$ pretrained on an in-distribution dataset $D_{\text{ID}}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, where elements of $D_{\text{ID}}$ are drawn from a joint distribution $P_{\mathcal{X}\mathcal{Y}}$, with support $\mathcal{X}\times\mathcal{Y}$. We denote with $P_{\text{ID}}$ its marginalization on $\mathcal{X}$. The OOD detection problem aims to determine whether an input sample originates from the in-distribution $P_{in}$ or not. Denoting with $\mathcal{Y}_{\text{OOD}}$ a set of labels such that $\mathcal{Y}\cap\mathcal{Y}_{\text{OOD}}=\emptyset$, OOD samples are drawn from a distribution $P_{\text{OOD}}$ which correspond to the marginalization on $\mathcal{X}$ of the joint distribution on $\mathcal{X}\times\mathcal{Y}_{\text{OOD}}$. i.e., they share the same input space $\mathcal{X}$ as in-distribution samples, but have labels outside $\mathcal{Y}$. Shifting from $P_{\text{ID}}$ to $P_{\text{OOD}}$ corresponds to a semantic change in label space.
@@ 3.1 Problem Setting
$d(\mathbf{x};s,f)=\begin{cases}\text{ID}&\text{if }s(\mathbf{x};f)\geq\lambda\\ \text{OOD}&\text{if }s(\mathbf{x};f)<\lambda\end{cases}$
where samples with high scores are classified as in-distribution, according to the threshold $\lambda$. For example, to compute the standard FPR95 metric [51], the threshold $\lambda$ is chosen such that it correctly classifies $95\%$ of ID held-out data. An ideal OOD score function should capture differences in model outputs between samples drawn from $P_{\text{ID}}$ and $P_{\text{OOD}}$, effectively, determining when the model encounters inputs from classes it was not trained on.
Algorithm 1 LAFO (Look Around and Find Out)
2:OOD score $s$
3:function LAFO($\mathbf{x},f,\bm{\mu}_{\text{ID}},\alpha$)
5:     $\mathbf{z}=f_{1}\circ\ldots\circ f_{L-1}(\mathbf{x})$triangleright Compute the penultimate last layer features $\mathbf{z}$
6:     $score\leftarrow-\infty$
8:         compute $\mathbf{z}_{db}$ as in Eq.1
9:         compute $\theta_{\hat{y},y^{\prime}}(\mathbf{z})$ using Eq. 2
10:         compute $\tilde{s}(\mathbf{z})$ using Eq. 3
11:         if $\tilde{s}(\mathbf{z})\geq score$ then
12:              $score=\tilde{s}(\mathbf{z})$
13:         end if
14:     end for
15:     return $score$triangleright Returns the maximum score across all other classes
16:end function
3.2 OOD Detection with Relative Angles
This section presents our OOD score function, which relates the feature representations with the decision boundaries using relative angles in the feature space to discriminate between ID and OOD samples. Figure 1 provides a geometric visualization of our method. Our approach leverages the geometric relationships between three key points in the feature space: (i) the initial representation of a sample, (ii) its projection onto the decision boundary, and (iii) the mean of in-distribution features.
We propose using the relation between feature representations and decision boundaries by deriving closed-form plane equations for the decision boundaries between any two classes. Specifically, we examine the angle formed between the feature representation vector and its projection onto the decision boundary. However, this angle is sensitive to the choice of origin, creating an ambiguity as the geometric relationship between the feature representation and the decision boundary should be translation-invariant. To address this, we propose to represent features in a reference frame relative to the mean of the in-distribution samples. Therefore, we incorporate in-distribution characteristics by centering around its mean, while ensuring scale and translation invariance.
We observe that the angle between the centered representation and its projection onto the decision boundary is larger for ID data, indicating them requiring higher cost to change their label which captures the model’s confidence. In contrast, for OOD data, angle is smaller since they are expected to be more unstable, as they do not contain strong clues about their predicted classes (see Figure 1 for a conceptual explanation and the ID/OOD histograms in Appendix A.1 for empirical evidence).
Our framework provides a concise scoring with useful properties such as translation and scale invariance. These properties enable LAFO to be used in conjunction with existing activation shaping algorithms and allow for confidence aggregation across different models through score summation.
@@ 3.3 Features on the Decision Boundary
where $\mathbf{w}_{y_{1}}$ (or $\mathbf{w}_{y_{2}}$) denotes the the row vectors of $\mathbf{W}$ corresponding to class ${y_{1}}$ (respectively ${y_{2}}$) and similarly, $b_{y_{1}},b_{y_{2}}$ are the bias values corresponding to classes $y_{1}$ and $y_{2}$. Intuitively, given a fixed classifier, this equation is satisfied for all $\mathbf{z}$’s such that their corresponding logits for class $y_{1}$ and $y_{2}$ are equal. Then, feature representations can be projected orthogonally onto the hyperplane that defines the decision boundary:
$\mathbf{z}_{db}=\mathbf{z}-\frac{(\mathbf{w}_{y_{1}}-\mathbf{w}_{y_{2}})^{T}\mathbf{z}+(b_{y_{1}}-b_{y_{2}})}{||\mathbf{w}_{y_{1}}-\mathbf{w}_{y_{2}}||^{2}}(\mathbf{w}_{y_{1}}-\mathbf{w}_{y_{2}})$ (1)
Let $\bm{\mu_{\text{ID}}}\in\mathbb{R}^{D}$ be the mean of the in-distribution feature representations. Centering w.r.t. $\bm{\mu_{\text{ID}}}$ corresponds to shifting the origin to $\mu_{\text{ID}}$. In this new reference frame, three key points form a triangle in $D$-dimensional space: the centered feature vector ($\mathbf{z}-\bm{\mu_{\text{ID}}}$), its projection onto the decision boundary ($\mathbf{z_{db}}-\bm{\mu_{\text{ID}}}$) and the new origin (see Figure 1). Then, rather than the absolute distance between $\mathbf{z}$ and $\mathbf{z}_{db}$, we consider the relative angle $\theta_{y_{1},y_{2}}(\mathbf{z})$ from the in-distribution feature representation’s reference frame: this captures the position of features and the decision boundaries with respect to the in-distribution data, while also being scale invariant:
$\theta_{y_{1},y_{2}}(\mathbf{z})=\arccos{\left(\frac{<\mathbf{z}-\bm{\mu_{\text{ID}}},\mathbf{z}_{db}-\bm{\mu_{\text{ID}}}>}{||\mathbf{z}-\bm{\mu_{\text{ID}}}||\cdot||\mathbf{z}_{db}-\bm{\mu_{\text{ID}}}||}\right)}$ (2)
Our score function extracts the maximum discrepancy of the relative angles between the centered feature representation and its projections on $DB_{\hat{y},y^{\prime}}$, where $\hat{y}=\operatorname*{arg\,max}_{y\in\mathcal{Y}}g(\mathbf{z})$ and $y^{\prime}\in\mathcal{Y}$, $y^{\prime}\neq y$. Therefore for a sample $\mathbf{x}\in\mathcal{X}$, given $\mathbf{z}=f_{1}\circ...\circ f_{L_{1}}(\mathbf{x})$ we can write the score $s(\mathbf{x},f)$ as a function of $\mathbf{z}$:
@@
Properties. Intuitively, our score function captures several key aspects:
@@
  • Maximum Discrepancy. By considering the maximum angle across all non-predicted classes, we identify the ’furthest’ relative decision boundary, extracting model’s confidence on least likely class that an instance might belong.
  • Scale Invariance. Unlike absolute distances, angles remain consistent even if the feature space is scaled, allowing for fair comparisons between different models.
Relation with the state-of-the-art fDBD  [25]). We now provide a geometric interpretation for the score function fDBD. Using our analysis, we identified that their score can directly be mapped into the triangle we formed in Figure 1. For a sample $\mathbf{x}\in\mathcal{X}$:
$\text{fDBD}(\mathbf{z})=\frac{d(\mathbf{z},\mathbf{z}_{db})}{||\mathbf{z}-\bm{\mu_{\text{ID}}}||_{2}}$
where $\mathbf{z}\in\mathbb{R}^{D}$ is the feature representations of the input $x\in\mathcal{X}$, $\mathbf{z}_{db}\in\mathbb{R}^{D}$ is its projection onto the decision boundary, and $d(\cdot,\cdot)$ is the euclidean distance. Although seemingly unrelated, we can connect this score to our relative angle and demonstrate that the regularization term on the denominator brings a term that does not effectively discriminate between OOD and ID. As a result, impeding fDBD’s performance.
Using translation invariance of the euclidean distance, the same score can be written as:

### v2

# Out-of-Distribution Detection with Relative Angles

@@ Abstract
Deep learning systems deployed in real-world applications often encounter data that is different from their in-distribution (ID). A reliable model should ideally abstain from making decisions in this out-of-distribution (OOD) setting. Existing state-of-the-art methods primarily focus on feature distances, such as k-th nearest neighbors and distances to decision boundaries, either overlooking or ineffectively using in-distribution statistics. In this work, we propose a novel angle-based metric for OOD detection that is computed relative to the in-distribution structure. We demonstrate that the angles between feature representations and decision boundaries, viewed from the mean of in-distribution features, serve as an effective discriminative factor between ID and OOD data. We evaluate our method on nine ImageNet-pretrained models. Our approach achieves the lowest FPR in 5 out of 9 ImageNet models, obtains the best average FPR overall, and consistently ranking among the top 3 across all evaluated models. Furthermore, we highlight the benefits of contrastive representations by showing strong performance with ResNet SCL and CLIP architectures. Finally, we demonstrate that the scale-invariant nature of our score enables an ensemble strategy via simple score summation. Code is available at https://github.com/berkerdemirel/ORA-OOD-Detection-with-Relative-Angles.
@@ 1 Introduction
A trustworthy deep learning system should not only produce accurate predictions, but also recognize when it is processing an unknown sample. The ability to identify when a sample deviates from the expected distribution, and potentially rejecting it, plays a crucial role especially in safety-critical applications, such as medical diagnosis  (Fernando et al., 2021), driverless cars  (Bogdoll et al., 2022) and surveillance systems  (Diehl and Hampshire, 2002). The out-of-distribution (OOD) detection problem addresses the challenge of distinguishing between in-distribution (ID) and OOD data – essentially, drawing a line between what the system knows and what it does not.
@@
Various approaches have been proposed for OOD detection, mainly falling into two categories: (i) methods that suggest model regularization during training  (Lee et al., 2018; Hendrycks et al., 2019; Meinke and Hein, 2020), and (ii) post-hoc methods, which leverage a pre-trained model to determine if a sample is OOD by designing appropriate score functions  (Peng et al., 2024; Hendrycks and Gimpel, 2022; Sun et al., 2022). Post-hoc methods are more advantageous for their efficiency and flexibility, as they can be applied to arbitrary pre-trained models without retraining. These approaches are often categorized based on the domain of their score functions, i.e., at which representational abstraction level they assess if a sample is OOD or not. Earlier techniques focus on measuring the model confidence in the logits space  (Hendrycks and Gimpel, 2022; Liu et al., 2020), while the recent works employ distance-based scores  (Sun et al., 2022; Sehwag et al., 2021) defined in the model feature space. While logit-based methods suffer from the overconfident predictions of neural networks  (Minderer et al., 2021; Lakshminarayanan et al., 2017; Guo et al., 2017), the recent success of distance-based techniques highlights that the relationships in the latent space can provide a richer analysis.
@@
A natural approach to feature representations is by checking their proximity to the decision boundaries  (Liu and Qin, 2024). Conceptually, this can be related to identifying hard-to-learn examples in data-efficient learning  (Joshi et al., 2024; Chen et al., 2023). OOD samples can be viewed as hard-to-learn since they do not share the same label distribution as ID data. The success of this approach has been directly showed in fDBD score from  Liu and Qin (2024). However, our derivations revealed that the regularization term they use to incorporate ID statistics introduces an additional term that does not correlate with ID/OOD separation, ultimately hindering their performance.
@@
In this work, we present OOD Detection with Relative Angles (ORA), a novel approach that exploits the relationship between feature representations and classifier decision boundaries, in the context of the mean statistics of ID features. Unlike the earlier techniques, ORA introduces a new angle-based measure that calculates the angles between the feature representations and their projection onto the decision boundaries, relative to the the mean statistics of ID features. Changing reference frame to the mean of ID features adds another layer of discriminatory information to the score, as it naturally incorporates the ID statistics to the distance notion, exploiting the disparity between ID and OOD statistics. Moreover, the scale-invariant nature of angle-based representations, as similarly observed in  Moschella et al. (2023), allows us to aggregate the confidence scores from multiple pre-trained models simply by summing their ORA scores. This enables to have a score that can be single model based or extended to ensemble of models. In summary, our key contributions include:
@@
  • We present a novel post-hoc OOD score, which computes the angles between the feature representation and its projection to the decision boundaries, relative to the mean of ID-features
  • We conduct an extensive evaluation on the ImageNet OOD benchmark using 9 model backbones, including modern transformer-based architectures and compare 12 detection methods. ORA achieves the best average FPR95 across all models, ranks in the top 3 for every model, and is the best-performing method on 5 out of 9 models.
  • We analyze the benefits of using contrastively learned features with ORA. Our method achieves the best performance on CIFAR-10 with ResNet18-SCL and on ImageNet with ResNet50-SCL, reducing average FPR95 by $0.88\%$ and $7.74\%$ respectively. We further validate this trend with the CLIP model, where ORA achieves strong results in both zero-shot ($25.85\%$ FPR95) and linear probing ($23.94\%$ FPR95) settings.
@@
  • ORA’s scale-invariance allows aggregation of confidence scores from multiple pre-trained models, enabling an effective ensemble. Our experiments show that the ORA ensemble reduces the FPR95 by $2.51\%$ on the ImageNet OOD benchmark compared to the best single model performance.
@@ 2 Related Work
Early methods addressing the OOD detection problem  Bevandić et al. (2019); Hendrycks et al. (2019) utilize additional datasets to represent out-of-distribution data, training models with both positive and negative samples. This approach assumes a specific nature of OOD data, potentially limiting its effectiveness when encountering OOD samples that deviate from this assumption during inference.  Malinin and Gales (2018) designed a network architecture to measure distributional uncertainty. In  Geifman and El-Yaniv (2019)’s work, they provide another architecture with an additional reject option to abstain from answering.  Ming et al. (2022); Du et al. (2022), focused on synthesizing outliers rather than relying on auxilary datasets. On the other hand,  Van Amersfoort et al. (2020); Wei et al. (2022) argued that overconfident predictions of the networks on OOD data are the problem to be mitigated. For example,  Van Amersfoort et al. (2020) puts an additional gradient penalty to limit the confidence of the network. Whereas,  Wei et al. (2022) tackled the same problem by enforcing a constant logit vector norm during training. Although it is natural to impose structures during training for better OOD detection, these methods face the trade-off between OOD separability and model performance. Moreover, such approaches lack the flexibility of post-hoc score functions, as they necessitate model retraining which can be both time-consuming and computationally expensive.
Score functions.
@@
Recently, developing score functions for pretrained models on ID data has gained attention due to its ease of implementation and flexibility. These methods typically either couple feature representations with distance metrics, or measure a model’s confidence using its logits. Beyond canonical works such as Maximum Softmax Probability (Hendrycks and Gimpel, 2022), ODIN score (Liang et al., 2018), Energy score (Liu et al., 2020), we observed many advancements in post-hoc score design. For example, the activation shaping algorithms such as ASH (Djurisic et al., 2023), Scale (Xu et al., 2024b), and ReAct (Sun et al., 2021), apply activation truncations to feature representations, reducing model’s confidence for OOD data. These methods can be used in conjunction with ORA improving the performance. Recent distance-based methods KNN (Sun et al., 2022), NNGuide (Park et al., 2023), R-Mah (Ren et al., 2021) and fDBD (Liu and Qin, 2024) successfully utilized the feature representations from networks. Both NNGuide and KNN assign a score to a sample based on the kth nearest neighbor in ID training set. R-Mah measures the relative Mahalanobis distance between class centroids and the overall data centroid. In contrast, fDBD assigns a score to a sample based on its estimate of the distance between the feature representation and the decision boundaries. Moreover, recent angle-based methods such as P-Cos (Galil et al., ) and R-Cos (Bitterwolf et al., 2023) assign scores based on the cosine similarity between representations and class centroids.
@@
Our work falls into the score function category, serving as a plug-in for any pre-trained model on ID data. ORA combines feature space and logit space methods by utilizing the relative angle between the feature representation and its projections to the decision boundaries. Among the existing works, the closest approach to our method is fDBD, which uses a lower bound estimate to the decision boundaries. However, the regularization term they introduced inadvertently includes a term in their equation that is uncorrelated with being OOD or ID and can change spuriously, impeding performance. In contrast, we provide a score function that effectively incorporates in-distribution context and maintains scale invariance, all without extra regularization terms.
@@
Figure 1: Geometric visualization of ORA for in-distribution (left) and out-of-distribution (right) cases. ORA focuses on the angular distance between the feature representation and the decision boundary, from the perspective of the in-distribution mean. The angle $\theta$ serves as the distinguishing factor between ID and OOD samples, with $\theta_{\text{ID}}>\theta_{\text{OOD}}$.
@@ 3 Method
Problem Setting. We consider a supervised classification setting with input space $\mathcal{X}$ and label space $\mathcal{Y}$, following the literature  Yang et al. (2024). Given a model $f:\mathcal{X}\rightarrow\mathbb{R}^{|\mathcal{Y}|}$ pretrained on an in-distribution dataset $D_{\text{ID}}=\{(\mathbf{x}_{i},y_{i})\}_{i=1}^{N}$, where elements of $D_{\text{ID}}$ are drawn from a joint distribution $P_{\mathcal{X}\mathcal{Y}}$, with support $\mathcal{X}\times\mathcal{Y}$. We denote its marginal distribution on $\mathcal{X}$ as $P_{\text{ID}}$. The OOD detection problem aims to determine whether an input sample originates from the in-distribution $P_{\text{ID}}$ or not. Let $\mathcal{Y}_{\text{OOD}}$ be a set of labels such that $\mathcal{Y}\cap\mathcal{Y}_{\text{OOD}}=\emptyset$. OOD samples are drawn from a distribution $P_{\text{OOD}}$ which is the marginal distribution on $\mathcal{X}$ of the joint distribution over $\mathcal{X}\times\mathcal{Y}_{\text{OOD}}$ i.e., they share the same input space $\mathcal{X}$ as in-distribution samples, but have labels outside $\mathcal{Y}$. Shifting from $P_{\text{ID}}$ to $P_{\text{OOD}}$ corresponds to a semantic change in the label space.
@@
[…]
where samples with high scores are classified as ID, according to the threshold $\lambda$. For example, to compute the standard FPR95 metric (Yang et al., 2024), the threshold $\lambda$ is chosen such that it correctly classifies $95\%$ of ID held-out data. An ideal OOD score function should capture differences in model outputs between samples drawn from $P_{\text{ID}}$ and $P_{\text{OOD}}$, effectively detecting inputs from unseen classes.
3.1 Out-of-Distribution Detection with Relative Angles
This section introduces our OOD score function, which leverages relative angles in feature space to relate feature representations to decision boundaries, distinguishing ID from OOD samples. Figure 1 provides a geometric visualization of our method. Our approach leverages the geometric relationships between three key points in the feature space: (i) the initial representation of a sample, (ii) its projection onto the decision boundary, and (iii) the mean of in-distribution features.
We propose using the relation between feature representations and decision boundaries by deriving closed-form plane equations for the decision boundaries between any two classes. Specifically, we examine the angle formed between the feature representation vector and its projection onto the decision boundary. However, this angle is sensitive to the choice of origin, creating an ambiguity as the geometric relationship between the feature representation and the decision boundary should be translation-invariant. To address this, we propose to represent features in a reference frame relative to the mean of the in-distribution samples. Therefore, we incorporate ID characteristics by centering around its mean, while ensuring scale and translation invariance.
We observe that the angle between the centered representation and its projection onto the decision boundary is larger for ID data, indicating them requiring higher cost to change their label which captures the model’s confidence. In contrast, for OOD data, angle is smaller since they are expected to be more unstable, as they do not contain strong clues about their predicted classes (see Figure 1 for a conceptual explanation).
Our framework provides a concise scoring with useful properties such as translation and scale invariance. These properties enable ORA to be used in conjunction with existing activation shaping algorithms and allow for confidence aggregation across different models through score summation.
@@ 3.2 Features on the Decision Boundary
where $\mathbf{w}_{y_{1}}$ (or $\mathbf{w}_{y_{2}}$) denotes the the row vectors of $\mathbf{W}$ corresponding to class $y_{1}$} (respectively $y_{2}$}) and similarly, $b_{y_{1}},b_{y_{2}}$ are the bias values corresponding to classes $y_{1}$ and $y_{2}$. Intuitively, given a fixed classifier, this equation is satisfied for all $\mathbf{z}$’s such that their corresponding logits for class $y_{1}$ and $y_{2}$ are equal. Then, feature representations can be projected onto the hyperplane that defines the decision boundary:
$\mathbf{z}_{db}=\mathbf{z}-\frac{(\mathbf{w}_{y_{1}}-\mathbf{w}_{y_{2}})^{T}\mathbf{z}+(b_{y_{1}}-b_{y_{2}})}{\|\mathbf{w}_{y_{1}}-\mathbf{w}_{y_{2}}\|^{2}}(\mathbf{w}_{y_{1}}-\mathbf{w}_{y_{2}})$ (1)
Let $\bm{\mu_{\text{ID}}}\in\mathbb{R}^{D}$ be the mean of the in-distribution feature representations. Centering w.r.t. $\bm{\mu_{\text{ID}}}$ corresponds to shifting the origin to $\mu_{\text{ID}}$. In this new reference frame, three key points form a triangle in $D$-dimensional space: the centered feature vector ($\mathbf{z}-\bm{\mu_{\text{ID}}}$), its projection onto the decision boundary ($\mathbf{z_{db}}-\bm{\mu_{\text{ID}}}$) and the new origin (see Figure 1). Then, rather than the absolute distance between $\mathbf{z}$ and $\mathbf{z}_{db}$, we use the relative angle $\theta_{y_{1},y_{2}}(\mathbf{z})$ from the in-distribution feature representation’s reference frame. This captures the position of features and the decision boundaries relative to the in-distribution data, while also being scale invariant:
$\theta_{y_{1},y_{2}}(\mathbf{z})=\arccos{\left(\frac{\langle\mathbf{z}-\bm{\mu_{\text{ID}}},\mathbf{z}_{db}-\bm{\mu_{\text{ID}}}\rangle}{\|\mathbf{z}-\bm{\mu_{\text{ID}}}\|\cdot\|\mathbf{z}_{db}-\bm{\mu_{\text{ID}}}\|}\right)}$ (2)
Our score function captures the maximum discrepancy of the relative angles between the centered feature representation and its projections on $DB_{\hat{y},y^{\prime}}$, where $\hat{y}=\operatorname*{arg\,max}_{y\in\mathcal{Y}}g(\mathbf{z})$ is the predicted class, and $^{\prime}\in\mathcal{Y}, yy$$^{\prime}\neq y$. Therefore for a sample $\mathbf{x}\in\mathcal{X}$, given $\mathbf{z}=f_{1}\circ...\circ f_{L_{1}}(\mathbf{x})$ we can write the score $s(\mathbf{x},f)$ as a function of $\mathbf{z}$:
@@
Intuitively, our score function captures several key aspects:
@@
  • Maximum discrepancy. It selects the furthest decision boundary by finding the maximum angle across classes. This captures the model’s confidence in the least likely class.
  • Scale invariance. Unlike absolute distances, angles remain consistent even if the feature space is scaled, allowing for fair comparisons between different models. See Appendix  J for a theoretical justification of angle-based scores.
Relation with the state-of-the-art fDBD  (Liu and Qin, 2024)). We now provide a geometric interpretation for the score function fDBD. Using our analysis, we identified that their score can directly be mapped into the triangle we formed in Figure 1. For a sample $\mathbf{x}\in\mathcal{X}$:
$\text{fDBD}(\mathbf{z})=\frac{d(\mathbf{z},\mathbf{z}_{db})}{\|\mathbf{z}-\bm{\mu_{\text{ID}}}\|}$
where $\mathbf{z}\in\mathbb{R}^{D}$ is the feature representations of the input $x\in\mathcal{X}$, $\mathbf{z}_{db}\in\mathbb{R}^{D}$ is its projection onto the decision boundary, and $d(\cdot,\cdot)$ is the euclidean distance. Although seemingly unrelated, we can connect this score to our relative angle and demonstrate that the regularization term on the denominator brings a term that does not effectively discriminate between OOD and ID. Using translation invariance of the euclidean distance, the same score can be written as: